P=25 P= 20 Qd= 50 Qd= 100

1. Compute for the demand coefficient of price b-value

2. Compute for the level of demand independent of price

3. Compute for the demand function


P=25 P=20 Qs=75 Qs=50

1. Compute for the supply coefficient of d-value

2. Compute for the level of supply independent of price

3. Compute for the supplied function


1. What is the equilibrium price? round to the nearest hundredths

2. What is the quantity demand value?round to the nearest tenths

3 What is the quantity supplied value?round to the nearest tenths



1
Expert's answer
2021-08-29T16:50:38-0400

a:

given

P=25 P= 20 Qd= 50 Qd= 100

"(P_1=25, Q_1= 50)(P_2= 20 , Q_2= 100)"

The demand function is given as

"Q=a-bP\\eq...1\\50=a-b(25)....." using point 1

"eq..2\\100=a-b(20)....." using point 2

from equation (1),"50=a+b(25)=a"

"100=50+25b-20b\\50=5b\\10=b"

"a=50+25(10)=300"



1)

demand coefficient of price b-value

b=-10

2)

level of demand independent of price

a=300

3)

demand function

"Qd=300-10P"



b:

given

P=25 P=20 Qs=75 Qs=50

"(P_1=25, Q_1=75)(P_2=20 ,Q_2=50)"

The supply function is given as

"Q=a+bP\\eq...1\\75=a
+b(25)....." using point 1

"eq..2\\50=a-b(20)....." using point 2

from equation (1),"75-25b=a"

"50=75-25b+20b\\25=5b\\5=b"

"a=75+25(5)=-50"

1)

supply coefficient of price d-value

d=5

2)

level of supply independent of price

=-50

3)

demand function

"Qs=-50+5P"


c:

"Qs=Qd\\-50+5p=300-10p\\300+50=5p+10p\\350=15p\\p=23.333\\Q=300-10(23.333)\\Q=66.67"

1)p=23.333

2)Qd=66.67 units

3)Qs=66.67 units